Spring Green company rents a building with a total of 100,000 square feet, which are evenly divided between two floors. The space on the first floor is considered twice as valuable as that on the second floor. The total monthly rent for the building is $30,000. How much of the monthly rental expense should be allocated to a department that occupies 10,000 square feet on the first floor?
Answer #1

Answer : Rental Expenses = $ 4,000

Total Square Feets on the First Floor = 100,000/2 = 50,000

Total Rent for 50,000 Feets = 30,000*2/3 = 20,000

Rent Allocated to the Department = 10,000/50,000*20,000 = $ 4,000 (Answer)


answered by: ANURANJAN SARSAM